OpenStudy (kaylala):

Trigonometric Equation tanA cosA + √(3) cosA = 0

OpenStudy (kaylala):

One more @hartnn

Parth (parthkohli):

tan(A) is sin(A)/cos(A).

Parth (parthkohli):

Sorry.

OpenStudy (kaylala):

No prob

OpenStudy (kaylala):

So it shall be sin A + cosA= 0 @hartnn ?

hartnn (hartnn):

and where did sqrt 3 go ?

OpenStudy (kaylala):

Oh insert that before cos A Sorry

OpenStudy (kaylala):

What's next?

hartnn (hartnn):

ok, divide both sides by 2

OpenStudy (kaylala):

1/2sinA + sqr(3) / 2

OpenStudy (kaylala):

Then?

hartnn (hartnn):

(1/2)sinA + (sqr(3) / 2) cos A = 0

hartnn (hartnn):

you know what is sin (x+y) =... ?

OpenStudy (kaylala):

Nope

OpenStudy (kaylala):

What's it?

Parth (parthkohli):

Bring what hartnn wrote in terms of sinA/cosA.

OpenStudy (kaylala):

You mean tan?

OpenStudy (kaylala):

I really don know

OpenStudy (kaylala):

*dont

Parth (parthkohli):

Exactly. Bring sinA/cosA on one side.

OpenStudy (kaylala):

How?

OpenStudy (kaylala):

I dont get you?

Parth (parthkohli):

1/2 sinA = -sqrt3/2 cosA 1/2 sinA/cosA = -sqrt3/2 sinA/cosA = -sqrt(3) tanA = -sqrt3

Parth (parthkohli):

tanA is -sqrt(3) for what A?

OpenStudy (kaylala):

Is that even possible? Moving to the other side?

Parth (parthkohli):

Yes. Why not? :D

OpenStudy (kaylala):

Okay. It's the 1st time i saw that. Haha

OpenStudy (kaylala):

What's next then?

Parth (parthkohli):

As I asked, for what value of A is tan(A) equal to -sqrt(3)?

OpenStudy (kaylala):

A=60,300?

OpenStudy (kaylala):

Oh that's - How will that be?

Parth (parthkohli):

Hmm.

OpenStudy (kaylala):

No solution?

Parth (parthkohli):

There is, actually. tan is negative when sin and cos have opposite signs. @hartnn Aliter solution?

hartnn (hartnn):

you just can't cancel cos A cos A (tan A + sqrt 3) = 0 cos A = 0 tan A = -sqrt 3 we need to solve these both

Parth (parthkohli):

90 deg is not a solution. tanA will be undefined. :-/ Can cancel.

OpenStudy (kaylala):

How's that?

hartnn (hartnn):

ofcourse, but thats true for this particular problem only. in general, we can't do this. we need to mention one solution as 90 and then discard it saying tan 90 is not defined. we need to show that analysis

Parth (parthkohli):

But yeah, the other solution is the same.

OpenStudy (kaylala):

What's gonna happen now?

Parth (parthkohli):

True, but we can write alongside that since cos cannot be zero, we can cancel.

hartnn (hartnn):

yes, thats my point, that we need to mention.

OpenStudy (kaylala):

Please elaborate later. Thanks. I really am lost and confused

Parth (parthkohli):

Kaylala, tanA + sqrt3 = 0.

OpenStudy (kaylala):

X=-sqrt3 Then what? @ParthKohli

OpenStudy (kaylala):

Please help @hartnn @ParthKohli

hartnn (hartnn):

ok, so u have problem finding angles for which tan A = -sqrt 3 ?

OpenStudy (kaylala):

Yea i do

OpenStudy (kaylala):

How to do this?

OpenStudy (kaylala):

@hartnn

hartnn (hartnn):

the ratio of sin A to cos A should give you -sqrt 3

OpenStudy (kaylala):

I havnt encountered such What will it be. @hartnn

hartnn (hartnn):

look for angles with sin A = sqrt 3/2 cos A = -1/2 or sin A = -sqrt 3/2 cos A = 1/2

OpenStudy (kaylala):

I see. Then? I dont know

OpenStudy (kaylala):

So i am to pict the bottom sin and cos?

hartnn (hartnn):

you will be able to find both the angles

OpenStudy (kaylala):

A={60,300, 30,330} ? @hartnn

OpenStudy (kaylala):

Am i right. ? @hartnn

OpenStudy (kaylala):

Brb

OpenStudy (anonymous):

This is what I got. Try to follow the steps.

hartnn (hartnn):

its only 120 and 300 the other method i was proposing, (1/2)sinA + (sqr(3) / 2) cos A = 0 let B = 60 sin 60 = sqrt 3/2 cos 60 = 1/2 cos 60 sin A + sin 60 cos A = 0 sin (A+60) = 0 A +60 = 0 A+60 = 180 A+60 = 360 A = -60 = 300 A = 180-60 = 120

OpenStudy (kaylala):

Okay i'm really confused which among the 3 answers and solutions is correct. @ParthKohli @hartnn @kurtsloane Could anyone help please. Thanks :)

OpenStudy (kaylala):

@Callisto Help. I'm confused with all these

OpenStudy (kaylala):

@phi @UnkleRhaukus @petiteme @Preetha

OpenStudy (unklerhaukus):

i see there are many methods

OpenStudy (unklerhaukus):

my favourite one is multiply by a half then use sum/difference trig formulas in reverse

OpenStudy (kaylala):

But whose answer is correct? @UnkleRhaukus In your opinion Who got it right

OpenStudy (kaylala):

Is: A= 60,300,30,330 Or A=60,300

OpenStudy (unklerhaukus):

well A=30° doesn't work,that would make LHS =2

OpenStudy (unklerhaukus):

and A=330° would make LHS=1

OpenStudy (kaylala):

Oh okay. Thanks

OpenStudy (unklerhaukus):

tan(300°)=-√3

hartnn (hartnn):

and tan 120 = -√3 so, A = {120,300}

OpenStudy (anonymous):

You know that tan is negative when either sin or cos are negative, but not both (because tan = sin/cos). So, since tan of your angle is -sqrt(3), you know that the angle is in either quadrant II or quadrant IV, because quadrant II x is negative and y is positive, and quadrant 4, x is positive and y is negative. You can use rules like that to help you figure out where the angle should be and quick-check your answers. But yeah A = {120,300}